Imagine, just imagine, I am in the mood to give away 7 extra credit points among my (only) 4 students. If all students have the same opportunity to get any of the points, what is the probability that each student gets at least one extra point?
Answer #1

Here let say for each student get x1 ,x2 ,x3 and x4 credit points, so here

x1 + x2 + x3 + x4 = 7

where

x1 ,x2 ,x3 ,x4 >= 0

So total number of possibilities here are as per combinamatrics formula = (7 + 4 -1)C(4-1) = 10C3 = 120

Now, we know that each of the student get atleast one extra point so here there is only three extra point are left for distribution, so if y1,y2 ,y3 and y4 are the extra credit received by the students.

So total number of possibilities here are as per combinamatrics formula = (3 + 4 -1)C(4-1) = 6C3 = 20

so here probability that atleast each student gets at least one extra point = 20/120 = 1/6